OpenStudy (sagewilson):

f(x) = 2x + 5, g(x) = X - 5 / 2 A. Graph the functions f(x) and g(x) on the same coordinate plane. You may use technology to create the graph of the functions or submit a handwritten graph. B. In two or more complete sentences, explain how to use the graphs of f(x) and g(x) to prove that the functions are inverses of each other. I don't understand what it is I have to do.

563blackghost (563blackghost):

First you would plot these two functions....Were you able to do this?

OpenStudy (sagewilson):

f(x) is blue g(x) is orange.

OpenStudy (sagewilson):

Would they be inverses of each other because they end up going in opposite directions of the coordinate plane?

563blackghost (563blackghost):

So so sorry Im trying to refresh on finding inverse....

OpenStudy (sagewilson):

It's okay.

OpenStudy (sagewilson):

You're helping me so there really is no need to apologize.

563blackghost (563blackghost):

You would need to use... \(\huge{f(g(x))=g(f(x))=x}\) So we would substitute x for the opposite equation.... So lets do `f(x)` first... We have \(\huge{f(x)=2x+5}\) Now we substitute `g(X)` for x.... \(\huge{f(x)=2(x-\frac{5}{2})+5}\) Simplify...

OpenStudy (sagewilson):

\[2x - 2\frac{ 5 }{ 2 } (im assuming) + 5\]

OpenStudy (sagewilson):

*im assuming

563blackghost (563blackghost):

I actually got \(\Large{f(x)=2x}\)

OpenStudy (sagewilson):

oh. How did you get that?

563blackghost (563blackghost):

Wait is it \(\Large{g(x)=\frac{x-5}{2}}\) or is it the way I have been writing?

OpenStudy (sagewilson):

it's \[g(x) = \frac{ x - 5 }{ 2 }\]

563blackghost (563blackghost):

OOOOOoooo well then I though it was the other way cause you didnt put parenthesis so I didnt know sorry

OpenStudy (sagewilson):

My fault, sorry.

563blackghost (563blackghost):

So our equation is.... \(\huge{f(x)=2(\frac{x-5}{2})+5}\) So 2 multiplied by a fraction of 2 cancels out the two... and 5 (fractions) added by 5 cancels so.... \(\huge{f(x)=x}\)

OpenStudy (sagewilson):

I would be left with f(x) = \[\frac{ x }{ 2 }\] right?

563blackghost (563blackghost):

Not quite.